# AHAMask: Reliable Task Specification for Large Audio Language Models without Instructions

### v1

@@ Abstract
Although current large audio language models (LALMs) extend text large language models (LLMs) with generic acoustic understanding abilities, they usually suffer from instruction sensitivity, where different instructions of the same intention can yield drastically different outcomes. In this work, we propose AHAMask, where we simply mask some of the attention heads in the decoder-only LLM backbone of LALMs, to trigger specific acoustic task functionalities without instructions. These masks are efficiently obtained by training on an LALM, with the number of trainable parameters equal to the attention head count in its LLM backbone. We show by experiments that applying such selective attention head masks achieves comparable or even better performance than using instructions, either on single or composite tasks. Besides achieving reliable acoustic task specification for LALMs, this also reveals that LALMs exhibit certain “functional pathways” in their attention heads.
@@ Introduction
The rise of large language models (LLMs) has brought a revolution in the speech and audio processing domain, resulting in a mature paradigm of large audio language models (LALMs) (Peng et al. 2024; Su et al. 2025; Yang, Ho, and Lee 2025; Arora et al. 2025). These LALMs extend text-based LLMs to “hear” and further understand audio input, typically with an audio encoder as the acoustic sensory receptor, and an LLM as the “brain” to process and response according to the audio input.
@@
However, the flexibility of natural language also brings the risk of instruction sensitivity. Even targeted at a specific task, LALMs can still exhibit a large degree of sensitivity to provided instructions, even when these instructions have identical task requirements but only differ in linguistic forms or template specifications. This sensitivity poses a significant challenge for LALMs, impacting both their ability to accurately follow instructions and maintain consistent task performance (Peng et al. 2024). This problem is also a consensus in the broader LLM research field. However, for LALMs, limited effort has been devoted to benchmarking instruction sensitivity, let alone addressing it.
Since instructions are expected control the functionality of LALMs, we choose to directly control the functionality inside the Transformer (Vaswani et al. 2017) structure of LALMs instead. This relates to the general explainability and functional partitioning of decoder-only Transformers, which have been particularly a trending research topic for text LLMs. Recently, Han et al. (2025) finds that simply masking certain attention heads in LLMs leads to specific task functionalities without instructions. This selective attention head masking mechanism does not apply any modification to the model parameters, revealing the fundamental existence of “functional pathway” among the attention heads. However, this discovery is limited to only the text modality instead of multi-modal LLMs like LALMs. Since audio understanding itself stands as a different functionality of perceiving physical signals, it is worth exploring whether there exist similar properties in LALMs.
Inspired by the previous works in text-based LLMs, we verify in this study that the “acoustic functional pathways” also exist in LALMs’ attention heads. Thus, we propose AHAMask (Acoustic Attention Head Mask) to achieve reliable task specification for LALMs, simply by masking some attention heads. Fig. 1 depicts the difference between a typical LALM with instructions and with only AHAMask. Importantly, the parameter count of our method is only equal to the number of attention heads in LALM’s decoder-only LLM backbone, which is even magnitudes smaller then previous parameter-efficient fine-tuning methods (1-2k parameters vs. millions of parameters). This means training an AHAMask is highly efficient. As these masks are binary in inference stage, the storage overhead of such masks is also neglectable, e.g. 200 bytes for SALMONN. AHAMask distinguishes itself from existing fine-tuning approaches by reducing the effective parameter count during inference, rather than maintaining or increasing it. We conduct experiments across famous LALMs and classic audio and speech understanding tasks, including single and composite multi-hop tasks. Our findings suggest that:
@@
This work considers LALMs, or also known as audio understanding models, as large Transformer-based models that can generate textual response to audio inputs and instructions, Popular LALMs in the current literature usually make use of an audio encoder and a pretrained backbone text LLM (see Fig. 1 left). The audio encoders are self-supervised or supervised learning models to provide high-dimensional contextual representations of input audio signals. For example, the encoder of Whisper (Radford et al. 2023) and BEATs (Chen et al. 2023) are used and compressed together by Q-former (Li et al. 2023) in SALMONN (Tang et al. 2024) for speech and general audio input, respectively. These extracted embeddings are fed to an LLM backbone, which is typically initialized from a pretrained text LLM and finetuned on audio data, such as by LoRA (Hu et al. 2022). After training, the backbone LLM is able to understand and answer questions on acoustic information in various perspectives, such as spoken content, prosody, speaker traits, and even audio events. Recently, rapid progress has been made in this direction (Hu et al. 2024; Lu et al. 2024; Ding et al. 2025; Geng et al. 2025; Goel et al. 2025; Sakshi et al. 2025), making LALMs a successful and necessary component in artificial general intelligence.
Instruction Following and Sensitivity in LLMs
The instruction sensitivity issue in LLMs usually comes in the form of instruction non-following. For text-only LLMs, several benchmarks (Zeng et al. 2024; Zhou et al. 2023; Qin et al. 2024; Lou, Zhang, and Yin 2024; Wen et al. 2024) for evaluating instruction-following capabilities have been established. Optimization methods for improving instruction-following ability include steering vectors (Stolfo et al. 2025; He et al. 2025), supervised finetuning or reinforcement learning with sophisticated data pipelines (Xu et al. 2024; Dong et al. 2025; An et al. 2025; Agrawal et al. 2025). Additionally, prompt sensitivity is widely observed (Chatterjee et al. 2024; Sclar et al. 2024; Cao et al. 2024; Zhuo et al. 2024; Razavi et al. 2025); for instance, minor modifications such as altering punctuation marks alone can degrade LLM performance significantly (Sclar et al. 2024). In contrast, efforts to benchmark instruction-following ability and sensitivity in LALMs are ongoing but comparatively limited (Lu, Kuan, and Lee 2025; Gao et al. 2025).
@@
In text-based LLMs, researchers have demonstrated that some Transformer modules exhibit certain functionality. For example, some Transformers layers can be swapped for cross lingual transfer (Bandarkar et al. 2025). It is also found that the neurons in feed forward networks (FFNs) store knowledge, such as language (Zeng et al. 2025), concepts (Rai and Yao 2024), and tasks (Xiao et al. 2024). Analysis on the attention heads also show that they encourage instruction following (Wu et al. 2024), transport a compact representation of tasks (Todd et al. 2024), mitigate knowledge conflicts (Jin et al. 2024), etc. Recently, Han et al. (2025) finds that learning-based attention head selection can form specific task functional pathways in LLMs without instructions. These findings indicate LLMs possibly have some internal modularity. Note that this also has some similarity with studies in functional partitioning of the brain from neuroscience (Bertolero, Yeo, and D’Esposito 2015; Wig 2017). To the best of our knowldge, such partitioning and modularity have not been explored in LALMs so far.
@@
Since the backbone LLMs act as “brains” in LALMs, we seek for acoustic functional pathways inside those LLMs. Modern LLMs are decoder-only Transformers (Vaswani et al. 2017), each of whose layer consists of multi-head attention (MHA), normalization, feed forward network and skip connections. Denote $\mathbf{X}\in\mathbb{R}^{l\times d}$ as the input sequence with length $l$ and dimension $d$, the attention operation at the $i$-th layer and $j$-th attention head is defined as
@@
where $\mathbf{W}_{Q}^{(i,j)},\mathbf{W}_{K}^{(i,j)},\mathbf{W}_{V}^{(i,j)}$ are query, key and value projections with shape $d\times d_{\text{head}}$, respectively. The output of an MHA module is then the weighted sum of all attention head outputs in that layer, i.e. $\operatorname{MHA}_{i}(\mathbf{X})=\sum_{j=1}^{h}\mathbf{Y}^{(i,j)}\mathbf{W}_{O}^{(i,j)}$ where $\mathbf{W}_{O}^{(i,j)}\in\mathbb{R}^{d_{\text{head}}\times d}$ is the output projection. These MHA modules are the most important part of a Transformer, and have been a long-standing research topic in LLMs (Vaswani et al. 2017; Lin et al. 2022; Zhao et al. 2023, 2024; Zheng et al. 2024).
@@
With this notation, we can treat $\mathcal{M}$ as the trainable parameters, and use Eq.(2) to train an AHAMask. As $\mathcal{M}$ is a discrete variable, we use Gumbel-Sigmoid (Jang, Gu, and Poole 2017; Geng et al. 2020; Han et al. 2025) for gradient estimation. Specifically, let $\mathbf{M}\in\mathbb{R}^{n\times h}$ be the mask logits of all heads. In the forward process, the discrete mask $\mathcal{M}$ is obtained by
@@
where $\sigma(\cdot)$ is the sigmoid function, $\mathbb{I}(\cdot)$ is the indicator function, and $\mathbf{G}\in\mathbb{R}^{n\times h}$ is Gumbel noise sampled by $g_{k}=-\log(-\log\mathbf{\epsilon}_{k}),\forall g_{k}\in\mathbf{G}$ using $\epsilon_{k}\overset{\mathrm{i.i.d}}{\sim}\operatorname{Uniform}(0,1)$. In the backward pass, gradients on $\mathcal{M}$ (the hard mask) is grafted to $\mathbf{S}$ (the soft mask), using straight-through estimator (Bengio, Léonard, and Courville 2013). Scalar $\tau>0$ is a temperature hyperparameter used to control the sharpness of Gumbel approximation. Therefore, we only train the mask logits $\mathbf{M}$ using gradient descent. In inference, we obtain the discrete mask as $\mathcal{M}=\mathbb{I}(\mathbf{M}\geq 0)$. Also, $\mathbf{M}$ assigns each attention head an importance weight, which can be utilized to explore the gradual formation of functional pathways. Typically, with approximately 3 to 15 billion parameters, LALMs contain only a few thousand attention heads within their LLM backbone. This means the number of trainable parameters in AHAMask remains minimal, and the training process is highly efficient.
@@
Figure 2: Instruction sensitivity experiments on SALMONN. Different colors and columns denote different types of variations.
@@
In this work, we conduct studies on three well-known open-sourced LALMs: SALMONN (Tang et al. 2024), Qwen2Audio-Instruct, and Qwen2Audio-Base (Chu et al. 2024). SALMONN incorporates Whisper and BEATs as acoustic encoders, and vicuna v1.1 (Chiang et al. 2023) 13B as its backbone LLM. Each of the two variants of Qwen2Audio use Whisper-style encoders and Qwen-7B (Bai et al. 2023) as their backbone LLM. Qwen2Audio-Instruct is a supervised fine-tuned outcome of Qwen2Audio-Base for better instruction following ability. All the three models have exhibited good audio understanding abilities in various tasks as general LALMs (Sakshi et al. 2025).
@@
To discover atomic acoustic functionalities, we consider the most well-defined classic audio understanding tasks in different aspects. These include automatic speech recognition (ASR) and gender recognition (GR) with LibriSpeech (Panayotov et al. 2015), speech emotion recognition (SER) with IEMOCAP (Busso et al. 2008), automatic speaker verification (ASV) with VoxCeleb1 (Nagrani, Chung, and Zisserman 2017), automatic audio captioning (AAC) with AudioCaps (Kim et al. 2019), speech to text translation (S2TT) with CoVoST2 (Wang, Wu, and Pino 2020), and overlapped speech recognition (OSR) with Libri2Mix (Cosentino et al. 2020). In GR, SER and ASV, LALMs are required to perform classification tasks , e.g. Male or Female in GR, and Yes or No in ASV. In ASV, two speech utterances are concatenated with a 0.1s silence as a single input. OSR requires LALMs to transcribe two overlapped speech utterances in any order. In S2TT and AAC, LALMs should apply more complex abilities to understand the audio input, and output translations or captions. These tasks span across purely linguistic (as ASR, OSR, S2TT), paralinguistic (SER, GR, ASV) to even general acoustic (AAC) understanding capabilities.
@@
Instruction Sensitivity in LALMs
It has been pointed out by Peng et al. (2024) that LALMs may exhibit non-negligible sensitivity to linguistic variations in the instructions. Following Peng et al. (2024) and Sclar et al. (2024), we present more evidence to showcase this issue. For each task, we first write a general instruction, and then apply the following variations:
@@
For Alter symbol, Semantically equal and Add constraint variations, we construct 5 different prompts in each type. Please refer to the appendix for detailed instructions. We take SALMONN as a typical LALM here. Notably, all 7 individual tasks here are comprehensively covered within the training data of SALMONN. Our preliminary experiments indicate that SALMONN demonstrates greater robustness to variations in instructions compared to Qwen2Audio. For decoding, we use deterministic beam search with beam size 4, so that the difference in model performance is only caused by different instructions.
The performance with instruction variations are visualized in Fig. 2. These results clearly indicate that SALMONN still has high sensitivity to instructions with the same intention. For example, in ASR task, changing the prompt to all capital letters will cause more hallucinations (repeating or generating upper case phonemes instead of words), thus raising the WER to 12%. In GR task, adding the constraint “Male or Female” makes the model almost always output “Male”. Similar phenomenon can also be observed with SER. For SALMONN, ASV and S2TT tasks exhibit better robustness to instruction variations. In almost all cases, using longer and more complex instructions is likely to cause a degradation in performance, sometimes even drastically as in ASR and SER. These results exemplify that instruction sensitivity remains a significant challenge in LALMs.
@@
We apply AHAMask on SALMONN and Qwen2Audio variants, on all the 7 single tasks. We compare LALM performance with our method (“w/ AHAMask”) against ① with clear instructions (w/ Instruction) from the “general” type in the previous section, ② without instructions (w/o Instruction), and ③ using a random mask with the same number of activated heads as AHAMask for each task. The detailed instructions we use for each LALM are listed in Table 7. For Qwen2Audio, since the two variants are identical in structure, we also evaluate ④ swapped AHAMask across variants.
@@
  • LALMs with only certain attention heads activated can achieve comparable or even better performance than using natural language instructions on all of the 7 single audio understanding tasks. Even though LALMs with natural language instructions sometimes still outperform those with only AHAMask, the difference in performance is very small (e.g. on ASV, S2TT and OSR tasks with SALMONN). Only on ASR and S2TT with Qwen2Audio-Base, which are the pretraining tasks of this model, using instructions achieves noticeably better performance compared to AHAMask.
@@
We present the results in Table 3. The three LALMs can barely generate answers in a specific format without clear instructions or with a random mask. In all the metrics, applying AHAMask achieves better performance than using instructions, especially for Qwen2Audio-Base where it is hard to instruct the model well. Also, LALMs can be sensitive to the ordering of sub-tasks in such composite tasks, e.g. the original SALMONN model can generate answers in the format of “GR|ASR” well, but fails in the opposite ordering. Even if the specific format is satisfied, the performance in each sub-task is not as good as when only doing single tasks. For SALMONN and Qwen2Audio-Instruct, the gap between different task orders is largely reduced using AHAMask, and the performance of each sub-task is closer to the single task scenarios. These results further indicate that AHAMask can control more complex behavior of LALMs in processing multiple aspects of acoustic information.
@@
Figure 4: Performance of SALMONN on 4 non-classification tasks with AHAMask in different percentage of activated attention heads. The orange triangle marker denotes the metric in Table 2, i.e. when probability threshold is 0.5. Note that WERs for ASR and OSR are visualized in logarithmic scale as there are extreme values (>>100) because of numerous repetitions.

### v3

@@ Abstract
Although current large audio language models (LALMs) extend text large language models (LLMs) with generic acoustic understanding abilities, they usually suffer from prompt sensitivity, where different instructions of the same intention can yield drastically different outcomes. In this work, we propose AHAMask, where we simply mask some of the attention heads in the decoder-only LLM backbone of LALMs, to trigger specific acoustic task functionalities without instructions. These masks are efficiently obtained by training on an LALM, with the number of trainable parameters equal to the attention head count in its LLM backbone. We show by experiments that applying such selective attention head masks achieves comparable or even better performance than using instructions, either on single or composite tasks. Besides achieving reliable acoustic task specification for LALMs, this also reveals that LALMs exhibit certain “functional pathways” in their attention heads.
@@ Introduction
The rise of large language models (LLMs) has brought a revolution in the speech and audio processing domain, resulting in a mature paradigm of large audio language models (LALMs) (peng2024survey; su2025audiosurvey; yang2025towards; arora2025landscape). In this work, we consider LALMs which are Transformer-based (vaswani2017attention) models that can generate textual response to audio inputs and instructions. They extend text-based LLMs to “hear” and further understand audio input, typically with an audio encoder as the acoustic sensory receptor, and an LLM as the “brain” to process and response according to the audio input.
@@
However, the flexibility of natural language also brings the risk of prompt sensitivity. Even targeted at a specific task, LALMs can still exhibit a large degree of sensitivity to provided instructions, even when these instructions have identical task requirements but only differ in linguistic forms or template specifications. This sensitivity poses a significant challenge for LALMs, impacting both their ability to accurately follow instructions and maintain consistent task performance (peng2024survey). This problem is also a consensus in the broader LLM research field. However, for LALMs, limited effort has been devoted to benchmarking prompt sensitivity, let alone addressing it.
Since instructions are expected to control the functionality of LALMs, we choose to directly control the functionality inside the Transformer (vaswani2017attention) structure of LALMs instead. This relates to the general explainability and functional partitioning of decoder-only Transformers, which have been a trending research topic for text LLMs. Recently, han2025heads finds that simply masking certain attention heads in LLMs leads to specific task functionalities without instructions. This selective attention head masking mechanism does not apply any modification to the model parameters, revealing the fundamental existence of “functional pathway” among the attention heads. However, this discovery is limited only to narrow applications in the text modality instead of multi-modal LLMs like LALMs. The acoustic functionalities in LALMs are fundamentally distinct from those of text LLMs, requiring a focus on multimodal alignment between continuous signals and discrete textual spaces, together with the comprehension of paralinguistic and non-linguistic information. Therefore, it is worth investigating whether similar properties exist in LALMs.
Inspired by the previous works in text-based LLMs, we apply the method in han2025heads and verify in this study that the “acoustic functional pathways” also exist in LALMs’ attention heads. Thus, we propose AHAMask (Acoustic Attention Head Mask) to achieve reliable task specification for LALMs, simply by masking some attention heads. Fig. 1 depicts the difference between a typical LALM with instructions and with only AHAMask. Importantly, the parameter count of our method is only equal to the number of attention heads in LALM’s decoder-only LLM backbone, which is even magnitudes smaller then previous parameter-efficient fine-tuning methods (1-2k parameters vs. millions of parameters). This means training an AHAMask is highly efficient. As these masks are binary in inference stage, the storage overhead of such masks is also neglectable, e.g. 200 bytes for SALMONN (tang2024salmonn). AHAMask distinguishes itself from existing fine-tuning approaches by reducing the effective parameter count during inference, rather than maintaining or increasing it. We conduct experiments across famous LALMs and classic audio and speech understanding tasks, including single and composite multi-hop tasks. Our findings suggest that:
@@
Popular LALMs in the current literature usually make use of an audio encoder and a pretrained backbone text LLM (see Fig. 1 left). The audio encoders are self-supervised or supervised learning models to provide high-dimensional contextual representations of input audio signals. For example, the encoder of Whisper (whisper) and BEATs (beats) are used and compressed together by Q-former (li2023blip) in SALMONN (tang2024salmonn) for speech and general audio input, respectively. These extracted embeddings are fed to an LLM backbone, which is typically initialized from a pretrained text LLM and finetuned on audio data, such as by LoRA (hu2022lora). After training, the backbone LLM is able to understand and answer questions on acoustic information in various perspectives, such as spoken content, prosody, speaker traits, and even audio events. Recently, rapid progress has been made in this direction (hu-etal-2024-wavllm; lu24c_interspeech_desta; ding2025kimi; geng2025osum; goel2025audio; sakshi2025mmau), making LALMs a successful and necessary component in artificial general intelligence.
LLM Instruction Following and Prompt Sensitivity
The prompt sensitivity issue in LLMs usually comes in the form of instruction non-following. For text-only LLMs, several benchmarks (zeng2024evaluating; zhou2023instruction; qin2024infobench; lou2024large; wen2024bench) for evaluating instruction-following capabilities have been established. Optimization methods for improving instruction-following ability include steering vectors (stolfo2025improving; he2025saif), supervised finetuning or reinforcement learning with sophisticated data pipelines (xu2024wizardlm; dong2025selfplay; an2025ultraifadvancinginstructionfollowing; agrawal2025enhancing). Additionally, prompt sensitivity is widely observed (chatterjee-etal-2024-posix; sclar2024quantifying; worstprompt_nips2024; zhuo-etal-2024-prosa; razavi2025benchmarking); for instance, minor modifications such as altering punctuation marks alone can degrade LLM performance significantly (sclar2024quantifying). In contrast, efforts to benchmark instruction-following ability and sensitivity in LALMs are ongoing but comparatively limited (lu2025speech; gao2025ifevalaudio).
@@
In text-based LLMs, researchers have demonstrated that some Transformer modules exhibit certain functionality. For example, some Transformers layers can be swapped for cross lingual transfer (bandarkar2025layer). It is also found that the neurons in feed forward networks (FFNs) store knowledge, such as language (zeng-etal-2025-converging), concepts (rai2024investigation), and tasks (xiao2024configurable). Analysis on the attention heads also show that they encourage instruction following (wu-etal-2024-language), transport a compact representation of tasks (todd2024function), mitigate knowledge conflicts (jin-etal-2024-cutting), etc. Recently, han2025heads finds that learning-based attention head selection can form specific task functional pathways in LLMs without instructions. These findings indicate LLMs possibly have some internal modularity. Note that this also has some similarity with studies in functional partitioning of the brain from neuroscience (bertolero2015modular; wig2017segregated). To the best of our knowldge, such partitioning and modularity have not been explored in LALMs so far.
@@
Figure 2: Prompt sensitivity experiments on SALMONN. Different colors and columns denote different types of variations.
Since the backbone LLMs act as “brains” in LALMs, we follow han2025heads and seek for acoustic functional pathways inside those LLMs. Modern LLMs are decoder-only Transformers, where each layer consists of multi-head attention (MHA), normalization, feed forward network and skip connections. Denote $\mathbf{X}\in\mathbb{R}^{l\times d}$ as the input sequence with length $l$ and dimension $d$, the attention operation at the $i$-th layer and $j$-th attention head is defined as
@@
where $\mathbf{W}_{Q}^{(i,j)},\mathbf{W}_{K}^{(i,j)},\mathbf{W}_{V}^{(i,j)}$ are query, key and value projections with shape $d\times d_{\text{head}}$, respectively. The output of an MHA module is then the weighted sum of all attention head outputs in that layer, i.e. $\operatorname{MHA}_{i}(\mathbf{X})=\sum_{j=1}^{h}\mathbf{Y}^{(i,j)}\mathbf{W}_{O}^{(i,j)}$ where $\mathbf{W}_{O}^{(i,j)}\in\mathbb{R}^{d_{\text{head}}\times d}$ is the output projection. These MHA modules are the most important part of a Transformer, and have been a long-standing research topic in LLMs (vaswani2017attention; lin2022survey; zhao2023survey; zhao2024explainability; zheng2024attention).
@@
With this notation, we can treat $\mathcal{M}$ as the trainable parameters, and use Eq.(2) to train an AHAMask. As $\mathcal{M}$ is a discrete variable, we use Gumbel-Sigmoid (jang2017categorical; geng-etal-2020-selective; han2025heads) for gradient estimation. Specifically, let $\mathbf{M}\in\mathbb{R}^{n\times h}$ be the mask logits of all heads. In the forward process, the discrete mask $\mathcal{M}$ is obtained by
@@
where $\sigma(\cdot)$ is the sigmoid function, $\mathbb{I}(\cdot)$ is the indicator function, and $\mathbf{G}\in\mathbb{R}^{n\times h}$ is Gumbel noise sampled by $g_{k}=-\log(-\log\mathbf{\epsilon}_{k}),\forall g_{k}\in\mathbf{G}$ using $\epsilon_{k}\overset{\mathrm{i.i.d}}{\sim}\operatorname{Uniform}(0,1)$. In the backward pass, gradients on $\mathcal{M}$ (the hard mask) is grafted to $\mathbf{S}$ (the soft mask), using straight-through estimator (bengio2013estimating). Scalar $\tau>0$ is a temperature hyperparameter used to control the sharpness of Gumbel approximation. Therefore, we only train the mask logits $\mathbf{M}$ using gradient descent. In inference, we obtain the discrete mask as $\mathcal{M}=\mathbb{I}(\mathbf{M}\geq 0)$. Also, $\mathbf{M}$ assigns each attention head an importance weight, which can be utilized to explore the gradual formation of functional pathways. Typically, with approximately 3 to 15 billion parameters, LALMs contain only a few thousand attention heads within their LLM backbone. This means the number of trainable parameters in AHAMask remains minimal, and the training process is highly efficient.
@@
In this work, we conduct studies on three well-known open-sourced LALMs: SALMONN (tang2024salmonn), Qwen2Audio-Instruct, and Qwen2Audio-Base (chu2024qwen2audio). SALMONN incorporates Whisper and BEATs as acoustic encoders, and vicuna v1.1 (vicuna2023) 13B as its backbone LLM. Each of the two variants of Qwen2Audio use Whisper-style encoders and Qwen-7B (qwen) as their backbone LLM. Qwen2Audio-Instruct is a supervised fine-tuned outcome of Qwen2Audio-Base for better instruction following ability. All the three models have exhibited good audio understanding abilities in various tasks as general LALMs (sakshi2025mmau).
@@
To discover atomic acoustic functionalities, we consider well-established benchmark tasks covering different aspects of audio understanding. These include automatic speech recognition (ASR) and gender recognition (GR) with LibriSpeech (librispeech), speech emotion recognition (SER) with IEMOCAP (iemocap), automatic speaker verification (ASV) with VoxCeleb1 (voxceleb), automatic audio captioning (AAC) with AudioCaps (kim2019audiocaps), speech to text translation (S2TT) with CoVoST2 (wang2020covost), and overlapped speech recognition (OSR) with Libri2Mix (cosentino2020librimix). In GR, SER and ASV, LALMs are required to perform classification tasks , e.g. Male or Female in GR, and Yes or No in ASV. In ASV, two speech utterances are concatenated with a 0.1s silence as a single input. OSR requires LALMs to transcribe two overlapped speech utterances in any order. In S2TT and AAC, LALMs should apply more complex abilities to understand the audio input, and output translations or captions. These tasks span across purely linguistic (as ASR, OSR, S2TT), paralinguistic (SER, GR, ASV) to even general acoustic (AAC) understanding capabilities.
@@
Prompt Sensitivity in LALMs
peng2024survey point out that speech LLMs exhibit non-negligible sensitivity to linguistic variations in the prompts. Following peng2024survey and sclar2024quantifying, we present more evidence to showcase this issue for general LALMs. For each task, we first write a general instruction, and then apply one of the following variations:
@@
For Alter symbol, Semantically equal and Add constraint variations, we construct 5 different prompts in each type. We take SALMONN as a typical LALM here. Notably, all 7 individual tasks here are comprehensively covered within the training data of SALMONN. Our preliminary experiments indicate that SALMONN demonstrates greater robustness to variations in instructions compared to Qwen2Audio. For decoding, we use deterministic beam search with beam size 4, so that the difference in model performance is only caused by different instructions.
The performance with instruction variations are visualized in Fig. 2. These results clearly indicate that SALMONN still has high sensitivity to instructions with the same intention. For example, in ASR task, changing the prompt to all capital letters will cause more hallucinations (repeating or generating upper case phonemes instead of words), thus raising the WER to 12%. In GR task, adding the constraint “Male or Female” makes the model almost always output “Male”. Similar phenomenon can also be observed with SER. For SALMONN, ASV and S2TT tasks exhibit better robustness to instruction variations. In almost all cases, using longer and more complex instructions is likely to cause a degradation in performance, sometimes even drastically as in ASR and SER. These results exemplify that prompt sensitivity remains a significant challenge in LALMs.
@@
We apply AHAMask on SALMONN and Qwen2Audio variants, on all the 7 single tasks. We compare LALM performance with our method (“w/ AHAMask”) against three other methods: ① with clear instructions (w/ Instruction) from the “general” type in the previous section, ② without instructions (w/o Instruction), and ③ using a random mask with the same number of activated heads as AHAMask for each task. For Qwen2Audio, since the two variants are identical in structure, we also evaluate ④ swapped AHAMask across variants.
@@
  • LALMs with only certain attention heads activated can achieve comparable or even better performance than using natural language instructions on all of the 7 single audio understanding tasks. Even though LALMs with natural language instructions sometimes still outperform those with only AHAMask, the difference in performance is very small (e.g. on ASV, S2TT and OSR tasks with SALMONN). Note that AHAMask is only supposed to reliably manifest specific functionalities instead of improving them, so these are still expected and supporting evidence. Only on ASR and S2TT with Qwen2Audio-Base, which are the pretraining tasks of this model, using instructions achieves noticeably better performance compared to AHAMask.
@@
We present the results in Table 3. In all the metrics, applying AHAMask achieves better performance than using instructions, especially for Qwen2Audio-Base where it is hard to instruct the model well. Also, LALMs can be sensitive to the ordering of sub-tasks in such composite tasks, e.g. the original SALMONN model can generate answers in the format of “GR|ASR” well, but fails in the opposite ordering. Even if the specific format is satisfied, the performance in each sub-task is not as good as when only doing single tasks. For SALMONN and Qwen2Audio-Instruct, the gap between different task orders is largely reduced using AHAMask, and the performance of each sub-task is closer to the single task scenarios. These results further indicate that AHAMask can control more complex behavior of LALMs in processing multiple aspects of acoustic information.
@@
Figure 4: Performance of SALMONN on 4 non-classification tasks with AHAMask in different percentage of activated attention heads. The orange triangle marker denotes the metric in Table 2